# FOLIO: Natural Language Reasoning with First-Order Logic

### v2

@@ Abstract
Simeng Han1 Hailey Schoelkopf1 Yilun Zhao1 Zhenting Qi2 Martin Riddell1 Wenfei Zhou1 James Coady1 David Peng1 Yujie Qiao1 Luke Benson1 Lucy Sun1 Alex Wardle-Solano1 Hannah Szabo1 Ekaterina Zubova1 Matthew Burtell1 Jonathan Fan3 Yixin Liu1 Brian Wong1 Malcolm Sailor1 Ansong Ni1 Linyong Nan1 Jungo Kasai4 Tao Yu4,5 Rui Zhang6 Alexander R. Fabbri8 Wojciech Kryściński8 Semih Yavuz8 Ye Liu8 Xi Victoria Lin7 Shafiq Joty8 Yingbo Zhou8 Caiming Xiong8 Rex Ying1 Arman Cohan1 Dragomir Radev1,8 1Yale University, 2Harvard University, 3Iowa City West High School 4University of Washington, 5University of Hong Kong 6Penn State University, 7Meta AI, 8Salesforce Research
@@ 1 Introduction
| FOLIO (ours) | 1,435 | FOL | Expert-written, Real-world | 4351 | 76 |
@@ 3 FOLIO Corpus Construction
We collected FOLIO through a carefully designed manual annotation process to achieve high-quality examples that necessitate complex logical reasoning. Writing natural language reasoning stories with FOL requires sufficient knowledge in both semantic parsing and first-order logic, as well as strong analytical skills. Given the complexities of such annotations, we selected annotators based on a few important criteria to ensures that our dataset is annotated with the highest level of precision and expertise, reflecting the complexity and nuance required for first-order logical reasoning. We include more details on annotator selection in Appendix A. We develop our dataset in six stages: WikiLogic collection, HybLogic collection, NL quality control, FOL quality control, NL-FOL alignment and FOL verification, spending 980 man-hours in total.
@@ 3.2 Quality control for NL sentences
Apart from grammar, we make sure the sentences in our dataset are highly natural. We also eliminate natural language ambiguity when it is possible. We include rules on eliminating ambiguity in Appendix D. Employing these rules effectively reduces the ambiguity of natural language in this reasoning dataset, but incurs the tradeoff of limiting variations in some usage of language. However, we note that there is still sufficient variation in terms of sentence structures and logical structures as shown in Table 1.
@@ 4.2 NL-FOL translation
Two metrics are adopted to evaluate NL-FOL translation to capture different aspects of the generation results: 1). Syntactic validity (SynV). The Syntactic Validity score measures whether the FOL formulas are syntactically valid. The score will be 1 if all FOL formulas of an example can pass the syntactic check and 0 otherwise 2). Inference Engine execution accuracy (EAcc). The group of translated FOL for premises and conclusions in one story is fed into our inference engine to output the truth value for each conclusion. We define the accuracy of the output labels as the execution accuracy. We leave for future work the design of a more reliable metric of NL-FOL translation.
@@ 5.3 Main results
We show that zero-shot prompting GPT-3.5 achieves better results than few-shot prompting text-davinci-002. Under few-shot NL prompting setting, LLaMA-13B achieves 33.63%, which is only slightly better than chance (33%). LLaMA-70B achieves 43.97%, around 10% better than LLaMA-13B. Text-davinci-002 achieves 49.53% and GPT-3.5 achieves 58.34%. GPT-4 achieves the best results among GPT series models.
@@
| Synv | EAcc | Sync | EAcc | |
@@ 6 Error Analysis
Models have higher accuracy on WikiLogic than on HybLogic As shown in Table 6, in logical reasoning, GPT-3.5 and GPT-4 achieve substantially lower results on HybLogic than on WikiLogic and the result is slightly higher than chance. We hypothesize that this is because HybLogic has high logical complexity that the SoTA LLMs like GPT-4 cannot solve yet. Since the examples in WikiLogic are created from scratch by humans, it is possible that LLMs have seen similar texts with similar logical patterns in the training data. However, fine-tuning RoBERTa-large achieves higher performance on HybLogic than on WikiLogic. This is likely because HybLogic is created from templates and some of the logical patterns can be learned during fine-tuning.
@@
Human evaluation on model outputs We conduct human evaluation on the GPT-4 model outputs with wrong truth value predictions. Approximately 65% of the time, the model struggles to construct accurate reasoning chains for complex problems with intricate steps, indicating a limited ability to solve problems with long reasoning chains. In 25% of cases, erroneous derivations occur within certain reasoning steps, highlighting potential inaccuracies and flaws in logical deductions. 5% of conclusions in FOLIO have a complex syntactic structure, posing comprehension challenges for GPT-4. 5% of outputs show that GPT-4 leverage commonsense reasoning to employ spurious shortcuts that lead to the wrong truth value for the conclusion.
@@ 6.1 Human performance
We collected truth value annotations of logical reasoning for FOLIO test set from expert and non-expert annotators. Our expert annotators are CS college students familiar with FOL. Non-expert annotators are community college or high school students who have not taken the SAT. Both expert and non-expert annotators are native English speakers. Expert annotations achieve an accuracy of while non-expert annotations achieves , with a gap of 34.16%. This shows that sufficient domain knowledge of FOL is necessary for good performance on FOLIO. The expert and GPT-4 gap is , suggesting significant room for model improvement.
@@ 8 Limitations
Due to the significant effort required for data annotation, we were unable further expand our dataset for investigating how the size of training data affects the performance of fine-tuning experiments. We encourage the community to apply our annotation protocol to expand this realistic and complex FOL reasoning story set.

### v3

@@ Abstract
Simeng Han1 Hailey Schoelkopf1 Yilun Zhao1 Zhenting Qi2 Martin Riddell1 Wenfei Zhou3 James Coady1 David Peng1 Yujie Qiao1 Luke Benson1 Lucy Sun1 Alex Wardle-Solano1 Hannah Szabo1 Ekaterina Zubova1 Matthew Burtell1 Jonathan Fan4 Yixin Liu1 Brian Wong1 Malcolm Sailor1 Ansong Ni1 Linyong Nan1 Jungo Kasai5 Tao Yu6 Rui Zhang7 Alexander R. Fabbri9 Wojciech Kryściński9 Semih Yavuz9 Ye Liu9 Xi Victoria Lin8 Shafiq Joty9 Yingbo Zhou9 Caiming Xiong9 Rex Ying1 Arman Cohan1 Dragomir Radev1,9 1Yale University, 2Harvard University, 3NVIDIA, 4Iowa City West High School 5University of Washington, 6University of Hong Kong 7Penn State University, 8Meta AI, 9Salesforce Research
@@ 1 Introduction
| FOLIO (ours) | 1,435 | FOL | Expert-written | 4351 | 76 |
@@ 3 FOLIO Corpus Construction
We collected FOLIO through a carefully designed manual annotation process to achieve high-quality examples that necessitate complex logical reasoning. Writing natural language reasoning stories with FOL requires sufficient knowledge in both semantic parsing and first-order logic, as well as strong analytical skills. Given the complexities of such annotations, we selected annotators based on a few important criteria to ensures that our dataset is annotated with the highest level of precision and expertise, reflecting the complexity and nuance required for first-order logical reasoning. 1). Our annotators are either college or graduate students who are native English speakers or possess near-native proficiency in English.4 2). They possess formal education in first-order logic, having either completed relevant coursework or undertaken self-directed studies in first-order logic or semantic parsing. At the NL quality check stage, only annotators who are experts in natural language processing or computational linguistics are involved. For the FOL quality check, only annotators who are experts in first-order logic are involved. We also give the annotators several training sessions on how to write a story, by providing them with detailed annotation guidelines. All stories and FOL annotations in FOLIO are written and reviewed by expert annotators, including CS undergraduate and graduate students, and senior researchers, who met the aforementioned criteria.
We develop our dataset in six stages: WikiLogic collection, HybLogic collection, NL quality control, FOL quality control, NL-FOL alignment and FOL verification, spending 980 man-hours in total.
@@ 3.2 Quality control for NL sentences
Apart from grammar, we make sure the sentences in our dataset are highly natural. All the sentences are first checked with a grammar checking tool, Grammarly. Our annotators who have graduated from or are senior students studying English Literature conducted a thorough round of review for grammatical correctness and language naturalness. We also eliminate natural language ambiguity when it is possible. We include rules on eliminating ambiguity in Appendix D. Employing these rules effectively reduces the ambiguity of natural language in this reasoning dataset, but incurs the tradeoff of limiting variations in some usage of language. However, we note that there is still sufficient variation in terms of sentence structures and logical structures as shown in Table 1.
@@ 4.2 NL-FOL translation
Two metrics are adopted to evaluate NL-FOL translation to capture different aspects of the generation results: 1). Syntactic validity (SynV). The Syntactic Validity score measures whether the FOL formulas are syntactically valid. The score will be 1 if all FOL formulas of an example can pass the syntactic check and 0 otherwise 2). Inference Engine execution accuracy (ExcAcc). The group of translated FOL for premises and conclusions in one story is fed into our inference engine to output the truth value for each conclusion. We define the accuracy of the output labels as the execution accuracy. We leave for future work the design of a more reliable metric of NL-FOL translation.
@@ 5.3 Main results
We show that zero-shot prompting GPT-3.5 achieves better results than few-shot prompting text-davinci-002. Under few-shot NL prompting setting, LLama-13B achieves 33.63%, which is only slightly better than chance (33%). LLama-70B achieves 43.97%, around 10% better than LLaMA-13B and obtains improvements of around 4% with Chain-of-thought prompting and Tree of Thought prompting. Text-davinci-002 achieves 49.53% and GPT-3.5 achieves 58.34%. GPT-4 achieves the best results among GPT series models.
@@
| Synv | ExcAcc | Sync | ExcAcc | |
@@ 6 Error Analysis
Models have higher accuracy on WikiLogic than on HybLogic As shown in Table 6, in logical reasoning, GPT-3.5 and GPT-4 achieve substantially lower results on HybLogic than on WikiLogic and the result is slightly higher than chance. We hypothesize that this is because HybLogic has high logical complexity that the SoTA LLMs like GPT-4 cannot solve yet while WikiLogic examples require shorter reasoning chains which the model is already capable of solving. Moreover, since the examples in WikiLogic are created from scratch by humans, it is possible that LLMs have seen similar texts with similar logical patterns in the training data. However, fine-tuning RoBERTa-large achieves higher performance on HybLogic than on WikiLogic. This is likely because HybLogic is created from templates and some of the logical patterns can be learned during fine-tuning.
@@
| Faulty path | 65% |
| Wrong derivation | 25% |
| Wrong syntactic comprehension | 5% |
| Spurious shortcut | 5% |
| Example |
| Premises 1. All children are human. 2. If someone is underage, then they are a child. 3. People are either underage or of age. 4. If someone is of age, then they can vote. 5. If someone is of age, they can legally get married. 6. If Jack is a child and a human, then Jack is neither able to vote nor able to get married. Conclusion -> Label: Jack is able to vote and get married. -> True. |
| GPT-4 Output |
| We know that all children are human (premise 1) and if someone is underage, they are a child (premise 2). People are either underage or of age (premise 3). If someone is of age, they can vote (premise 4) and get married (premise 5). If Jack is a child and a human, then Jack is neither able to vote nor get married (premise 6). We don’t have any information about Jack’s age, so we cannot determine if he is a child or of age. Therefore, we cannot determine if Jack is able to vote and get married. |
Human evaluation on model outputs We conduct human evaluation on the GPT-4 model outputs with wrong truth value predictions. As shown in Table 7, approximately 65% of the time, the model struggles to construct accurate reasoning chains for complex problems with intricate steps, leading to faulty reasoning paths and indicating a limited ability to solve problems with long reasoning chains. In 25% of cases, erroneous derivations occur within certain reasoning steps, highlighting potential inaccuracies and flaws in logical deductions. 5% of conclusions in FOLIO have a complex syntactic structure, posing comprehension challenges for GPT-4. 5% of outputs show that GPT-4 leverage commonsense reasoning to employ spurious shortcuts that lead to the wrong truth value for the conclusion. We provide a case study for the "Faulty path" scenario in Table 8. In this instance, the model can perform simple derivations from the premises, like "If someone is of age, they can vote and get married." However, because of the problem’s complexity, the model struggles to identify the essential intermediate steps and cannot ascertain the truth value of conclusions, such as "Jack is not a child."
@@ 6.1 Human performance
We collected truth value annotations of logical reasoning for FOLIO test set from expert and non-expert annotators. Our expert annotators are computer science college students familiar with FOL. Non-expert annotators are community college or high school students who have not taken the SAT. Both expert and non-expert annotators are native English speakers. Expert annotations achieve an accuracy of while non-expert annotations achieves , with a gap of 34.16%. This shows that sufficient domain knowledge of FOL is necessary for good performance on FOLIO. The expert and GPT-4 gap is , suggesting significant room for model improvement.
@@ 8 Limitations
We focus on collecting a very high-quality dataset in evaluating logical reasoning rather than merely a large dataset. Optimizing for quality required us to adopt a rigorous annotation process with domain experts selected based on a few important criteria as mentioned in Appendix A: Annotator Selection. Significantly scaling up this process would have required resources beyond our current means and we are unable further expand our dataset for investigating how the size of training data affects the performance of fine-tuning experiments. We encourage the community to apply our annotation protocol to expand this realistic and complex FOL reasoning story set.